# Characterizing Density and Gravitational Potential Fluctuations of the Interstellar Medium

### v1

@@ 4.2 Potential Fluctuation Vertical Structure
$(\delta\phi_{\mathbf{k}})_{\mathrm{m}}(z,t)=-\phi_{0}(t)\frac{\delta_{\mathbf{k}}(t)}{\tilde{k}}\chi_{\mathrm{m}}(\tilde{k},\tilde{z}),$ (49)
@@
We will use the exact expression in Section 4.2 when evaluating our models throughout the remainder of this paper, but for convenient reference, an approximate analytic expression with the correct limiting behavior is
$\chi_{\mathrm{m}}(\tilde{k},\tilde{z})\approx$ $(1-w)\alpha_{\mathrm{s}}\frac{\mathrm{sech}^{2}(\tilde{z}/\alpha_{\mathrm{s}})}{\left(\mathrm{sech}^{2p}(\tilde{z}/\alpha_{\mathrm{s}})+(\alpha_{\mathrm{s}}\tilde{k})^{p}\right)^{1/p}}$
@@
where the parameter $p>0$ controls the transition between the limits. Setting $p=3/2$ produces an approximation that is accurate to within $1\%$ when $\tilde{z}=0$ and $20\%$ across all relevant model parameters and $(\tilde{k},\tilde{z})$ values; setting $p=1$ instead yields errors of at most $12\%$ when $\tilde{z}=0$ and at most $25\%$ over the relevant parameters. Choosing $p=1$ for simplicity, the potential fluctuations within the disk ($|\tilde{z}|\ll 1$) are then given approximately by
$(\delta\phi_{\mathbf{k}})_{\mathrm{m}}\approx-\phi_{0}(t)\frac{\delta_{\mathbf{k}}(t)}{\tilde{k}}\left(\frac{(1-w)\alpha_{\mathrm{s}}}{1+\alpha_{\mathrm{s}}\tilde{k}}+\frac{w\alpha_{\mathrm{e}}}{1+\alpha_{\mathrm{e}}\tilde{k}}\right).$ (53)
Overall, our model allows us to relate the power spectrum of potential fluctuations to the power spectrum of linear surface density fluctuations we studied in Section 3.2,
@@
and we anticipate that the quantity $(P_{\delta\phi})_{\mathrm{m}}/\phi_{0}^{2}$ will have the same temporal dependence as the surface density fluctuations, as discussed in Section 3.3.
@@
Figure 12: The $k=0$ component of the potential fluctuations as a function of $\tilde{z}$ (top row) as well as selected $k>0$ components as a function of $\tilde{z}$ at fixed $\tilde{k}=\{2,4,6\}$ (bottom row, with $\tilde{k}$ increasing from top to bottom in each panel) for the R8 (left, purple) and LGR4 (right, orange) simulations. For the $k>0$ components, we show the square root of the power spectrum of potential fluctuations, which is the root-mean-square of each $\delta\phi_{\mathbf{k}}$. We normalize both rows by $\phi_{0}$; see Section 4.1 (and note the different vertical axis limits). For clarity, we only show the standard deviation (shaded region) around the time-averaged quantity (solid curve) for one value of $\tilde{k}$, but the fractional variations are comparable across all the curves. In each panel, the black dashed curve shows the result assuming a separable density with the fitted vertical profile shown in Figure 11, which leads to the vertical models given in Section 4.1 and Equation 54.
@@
The bottom row of Figure 12 shows the measured root-mean-square of the potential fluctuation Fourier amplitudes, $\sqrt{P_{\delta\phi}}$ (bottom row), normalized by $\phi_{0}$, as functions of $\tilde{z}$ for the R8 (purple, left column) and LGR4 (orange, right column) simulations. For the potential fluctuations, we show slices at $\tilde{k}=2$, $\tilde{k}=4$, and $\tilde{k}=6$ (top, middle, and bottom curves respectively); these correspond approximately to $kL\sim$ 10, 20, and 30 and were chosen as representative values capturing the first $\sim 1$ dex of surface density fluctuation power (see Figure 7). We show the time-averaged values as solid curves, and for clarity we only indicate the ±$\pm 1\sigma$ deviations around the average as shaded regions around one $\tilde{k}$ value; the fractional variations are similar for each $\tilde{k}$. The relatively small fractional variations about the mean for the R8 simulation indicate that much of the time-dependence is contained in the evolution of the gas disk’s center of mass and thickness; as in Figure 11, the LGR4 simulation exhibits larger deviations from the mean. In both simulations, comparing the mean to the fluctuations, we note that the typical amplitude of each fluctuation component is significantly less than the amplitude of the mean for all $|\tilde{z}|\gtrsim 1$. This indicates that while the potential fluctuations are extremely dynamically important near the midplane (where they can be key drivers of the evolution of vertical structure in the stellar disk), they may be negligible beyond one scale height away from it.
@@
Overplotted as black dashed curves in all panels are the predictions of the model, given by Section 4.1 for the mean and Equation 54 for the fluctuations. As expected given the excellent fit to the mean vertical profile shown in Figure 11, the model matches the measured mean vertical potential quite accurately. For each $\tilde{k}$ value of the fluctuations, the model curve typically falls within the ±$\pm 1\sigma$ range around the measured time-average curve, although for the smallest $\tilde{k}$ values, the power is somewhat overpredicted at $|\tilde{z}|\gtrsim 1$, and for the larger $\tilde{k}$ values, the power is somewhat underpredicted for $|\tilde{z}|\lesssim 1$. These discrepancies can be attributed to the fact that the underlying gas volume density fluctuations are more concentrated near the midplane than the mean vertical density profile, because the feedback is strongest there — we have so far been assuming that the fluctuations and mean profiles share the same vertical structure, neglecting this effect. We describe an extension to our model allowing for the fluctuation vertical profile to differ from the mean in Appendix B. Nevertheless, we find that even this simple model with a single function $\zeta_{\mathrm{m}}$ for both the $k=0$ and $k>0$ components captures many of the essential features of the spectrum of potential fluctuations.
@@
Figure 13: The square root of the power spectrum of potential fluctuations as a function of $\tilde{k}$ at fixed $\tilde{z}=\{0,1,1.5\}$ (increasing from top to bottom in each panel) for the R8 (left, purple) and LGR4 (right, orange) simulations. We normalize by $\phi_{0}$; see Section 4.1. This is the same quantity shown in panels (c) and (d) of Figure 12, except now as a function of $\tilde{k}$ instead of $\tilde{z}$. For clarity, we only show the standard deviation (shaded region) around the time-averaged quantity (solid curve) for one value of $\tilde{k}$, but the fractional variations are comparable across all the curves. In each panel, the black dashed curve shows the result assuming a separable density profile with the fitted vertical profile shown in Figure 11, according to Equation 54. The dotted curves indicate power laws of slope indicated by the text in each panel for reference.
@@
Figure 13 shows the same root-mean-square potential fluctuation amplitudes as the bottom row of Figure 12, but this time as a function of $\tilde{k}$ for the midplane $\tilde{z}=0$ (top curve in each panel), as well as $\tilde{z}=1$ and $\tilde{z}=1.5$ (middle and bottom curves respectively). Once again, we show the ±$\pm 1\sigma$ range about the time-average quantity for just one curve for clarity, but the fractional variations about the mean are comparable for all $\tilde{z}$ values. We again overplot the predictions from the model given by Equation 54 as black dashed curves. We indicate the approximate power law scalings measured from the simulation at low $\tilde{k}$ on each panel and plot dotted lines of the corresponding slope for comparison.
For the R8 simulation, the model is accurate across a wide range of $\tilde{k}$ for $|\tilde{z}|\geq 1$, and is reasonably accurate at the midplane for small values of $\tilde{k}$ as well, though the model amplitude drops off much more steeply with $\tilde{k}$ than the simulation. For the LGR4 simulation, the models are reasonably accurate at smaller $\tilde{k}$ values, but once again decay much more rapidly with $\tilde{k}$ than the simulation curve. From Equation 54 and the approximation presented in Section 4.2, we anticipate that at low $|\tilde{z}|$, the amplitude steepens from $\propto k^{-(n_{\delta}/2+1)}\approx k^{-2.1}$ to $\propto k^{-(n_{\delta}/2+2)}\approx k^{-3.1}$ as $\tilde{k}$ increases, while at larger $|\tilde{z}|$, the amplitude is $\propto k^{-(n_{\delta}/2+2)}\approx k^{-3.1}$ across all $\tilde{k}$ values. This steepening at low $|\tilde{z}|$ and the lack thereof at larger $|\tilde{z}|$ is apparent in both the R8 and LGR4 simulations, and in each case the model slopes are approximately consistent with (though somewhat steeper than) the simulations. The origin of the difference in slope between the model and the simulation at the midplane is the approximation that the surface density power is representative of the midplane density: in reality, a slice of the midplane density has a slightly shallower power spectrum than $P_{\delta}$, because small-scale features generated by turbulence and feedback at the midplane are smoothed over when integrated vertically to produce a surface density map. Note that the extension allowing the fluctuations to be more concentrated around the midplane described in Appendix B does improve the match to the curves over a somewhat wider range of $\tilde{k}$ values, though the difference at large $\tilde{k}$ persists. However, because the vast majority of the power is contained in the fluctuations at small $\tilde{k}$ (note the larger dynamic range shown on the vertical axis compared to Figure 12) where the absolute difference in fluctuation amplitude is not yet very large, our model still yields accurate results.
@@
Thus far, we have been comparing our model to the simulation outputs by studying only the amplitudes of each Fourier component of the fluctuations. However, another consequence of Equation 49 is that the phase of $(\delta\phi_{\mathbf{k}})_{\mathrm{m}}$ is entirely set by the phase of $\delta_{\mathbf{k}}$, since all the other factors are real. One simple way to assess whether the separable model is adequate in this regard is to reconstruct a potential fluctuation field from the measured $\delta_{\mathbf{k}}$ at a given snapshot (by summing up the components given by Equation 49 using Section 2) and comparing to the actual field at that time. If the phases of $\delta_{\mathbf{k}}$ and $\delta\phi_{\mathbf{k}}$ are well-correlated, then the reconstruction at each value of $z$ should have a similar morphology to the real field, but if the phases are not well-correlated, it may look entirely different.
@@
Ultimately, we have demonstrated that the simple separable model described above gives a good description of the potential fluctations as a function of both space and time. Together with the results regarding the linear surface density fluctuations presented in Section 3, Equation 54 provides a direct input for stellar dynamical studies of e.g., diffusion processes driven by external perturbations (J. Binney & C. Lacey, 1988), which we will explore in a future work (Modak et al. in prep).
5 Discussion
In this work, we have characterized the statistics of the (2D) surface density fluctuation fields of the ISM, and subsequently demonstrated that the structures of the (3D) volume density and gravitational potential fields are well-approximated by a separable model in which the 3D fluctuation fields essentially inherit the statistical structure of the 2D field (see Equation 38 and Equation 39), with a specified vertical profile (see Equation 44). Thus, our results can be summarized using our characterization of the surface density fluctuation spatio-temporal spectra in Section 3.3. With a realization of the surface density fluctuation $\delta(\mathbf{x},t)\equiv\Sigma(\mathbf{x},t)/\overline{\Sigma}(t)-1$ in hand, realizations of the corresponding volume density, vertical gravitational potential, and gravitational potential fluctuations can then be produced using Equation 38, Section 4.1, and Equation 49 respectively, and center-of-mass motion (Equation 36) or thickness variation (Equation 37) can be specified as desired, calibrated to the values reported in Table 4.
Physically, the content of Section 3.3 is that
@@ 5 Discussion
    fluctuation power is distributed across these modes as a steep power law with index $n_{\delta}\approx-2.3$ for linear surface density fluctuations $\delta$ (or $n_{s}\approx-2.8$ for logarithmic surface density fluctuations $s\equiv\ln(1+\delta)$), so that a majority of the power is concentrated at the largest spatial scales.
Below, we speculate on the implications and possible origins of the power-law spatial structure (Section 5.1), discuss how our results may be compared with observations (Section 5.2), and finally comment on caveats and possible generalizations of our characterization (Section 5.3).
5.1 Implications and Origins of Power-Law Spectra
An immediate consequence of the power-law spatial spectra of the surface density field is that ISM substructures are not well-described as compact “clouds” with finite size and mass: such an ISM would generically result in surface density power spectra with a peak at an intermediate scale corresponding to a typical cloud radius. In contrast to these classical models, the spectra we measure are relatively featureless, suggestive of a fractal, hierarchical structure. As an example of an application where this difference leads to qualitatively dissimilar results, we note that in the stellar-dynamical context, where it is natural to classify fluctuations by their scale (e.g., C. Hamilton et al. 2024), the large-scale ISM fluctuations that carry the most of the power drive scattering processes distinct from those of small-scale clouds. We will describe the implications of this fluctuation structure for stellar orbital heating and migration processes in Galactic disks in a future work (Modak et al. in prep).
@@ 5.1 Implications and Origins of Power-Law Spectra
Most analyses of turbulence in simulations, whether for highly idealized setups (e.g., isothermal, with driving in $\mathbf{k}$-space at large scales) or for setups that include more of the physics of the multiphase ISM, with injection of energy by supernovae (e.g., P. Padoan et al. 2016; J. R. Beattie et al. 2025), focus on velocity power spectra. Solenoidal components of the velocity have power spectra similar to the theoretical predictions of P. S. Iroshnikov (1964); R. H. Kraichnan (1965) or A. Kolmogorov (1941), while compressive components follow the steeper J. Burgers (1948) spectrum. While the perturbations in density are driven by the compressive components of the velocity field, the density and velocity spectra are expected to match only for pure acoustic waves. For very subsonic turbulent flows, simulations have recovered density spectra with $4\pi k^{2}P_{\mathrm{3D},\delta}(k)\propto k^{-2.2}$ (e.g. T. Gotoh & D. Fukayama, 2001; L. Pan et al., 2016), close to the expected index $-7/3$ for the pressure spectrum in incompressible turbulence (G. K. Batchelor, 1951) [^10] We have explicitly written $P_{\mathrm{3D}}$ to denote power spectra measured in 3D (and averaged over angles for $k=(k_{x}^{2}+k_{y}^{2}+k_{z}^{2})^{1/2}$), as distinct from our 2D power spectra. We note that for an isotropic spectrum, the projection in 2D simply selects $k_{z}=0$, with $P(k_{x},k_{y})\propto P_{\mathrm{3D}}(k=(k_{x}^{2}+k_{y}^{2})^{1/2})$ up to a normalization factor.. At higher Mach number $\cal M$, the density spectrum in simulations becomes flatter, approaching $4\pi k^{2}P_{\mathrm{3D},\delta}\propto k^{-1}$ at ${\cal M}\sim 10$ (e.g. J. Kim & D. Ryu, 2005; A. G. Kritsuk et al., 2007; L. Konstandin et al., 2016; L. Pan et al., 2016). Thus, for isotropic turbulence in a single-phase medium, the index of the 3D power spectrum is in the range $\sim-4$ to $-3$.
The spectra we find, $P_{\delta}\propto k^{-2.3}$ to $k^{-2.2}$ are considerably flatter than would be expected if we were simply projecting from the isotropic, single-phase models described above. The difference likely has more to do with the multiphase character of the gas than the fact that we are considering turbulence in a disk, since the ratio between the vertical dimension and the horizontal dimension in the simulation is modest, $2H/L\sim 0.6$. However, the cold gas is collected in small clouds with densities two orders of magnitude larger than those of the warm diffuse gas, while the density in superbubbles that fill a large fraction of the domain is two or more orders of magnitude below that of the warm gas.
5.2 Comparison with Observations

### v2

@@ IV.2 Potential Fluctuation Vertical Structure
[…]
@@
We will use the exact expression in Section IV.2 when evaluating our models throughout the remainder of this paper, but for convenient reference, an approximate analytic expression with the correct limiting behavior is
[…]
@@
[…]
Overall, our model allows us to relate the power spectrum of potential fluctuations to the power spectrum of linear surface density fluctuations we studied in Section III.2,
@@
and we anticipate that the quantity $(P_{\delta\phi})_{\mathrm{m}}/\phi_{0}^{2}$ will have the same temporal dependence as the surface density fluctuations, as discussed in Section III.3.
@@
Figure 12: The $k=0$ component of the potential fluctuations as a function of $\tilde{z}$ (top row) as well as selected $k>0$ components as a function of $\tilde{z}$ at fixed $\tilde{k}=\{2,4,6\}$ (bottom row, with $\tilde{k}$ increasing from top to bottom in each panel) for the R8 (left, purple) and LGR4 (right, orange) simulations. For the $k>0$ components, we show the square root of the power spectrum of potential fluctuations, which is the root-mean-square of each $\delta\phi_{\mathbf{k}}$. We normalize both rows by $\phi_{0}$; see Section IV.1 (and note the different vertical axis limits). For clarity, we only show the standard deviation (shaded region) around the time-averaged quantity (solid curve) for one value of $\tilde{k}$, but the fractional variations are comparable across all the curves. In each panel, the black dashed curve shows the result assuming a separable density with the fitted vertical profile shown in Figure 11, which leads to the vertical models given in Section IV.1 and Equation 54.
@@
The bottom row of Figure 12 shows the measured root-mean-square of the potential fluctuation Fourier amplitudes, $\sqrt{P_{\delta\phi}}$ (bottom row), normalized by $\phi_{0}$, as functions of $\tilde{z}$ for the R8 (purple, left column) and LGR4 (orange, right column) simulations. For the potential fluctuations, we show slices at $\tilde{k}=2$, $\tilde{k}=4$, and $\tilde{k}=6$ (top, middle, and bottom curves respectively); these correspond approximately to $kL\sim$ 10, 20, and 30 and were chosen as representative values capturing the first $\sim 1$ dex of surface density fluctuation power (see Figure 7). We show the time-averaged values as solid curves, and for clarity we only indicate the $\pm 1\sigma$ deviations around the average as shaded regions around one $\tilde{k}$ value; the fractional variations are similar for each $\tilde{k}$. The relatively small fractional variations about the mean for the R8 simulation indicate that much of the time-dependence is contained in the evolution of the gas disk’s center of mass and thickness; as in Figure 11, the LGR4 simulation exhibits larger deviations from the mean. In both simulations, comparing the mean to the fluctuations, we note that the typical amplitude of each fluctuation component is significantly less than the amplitude of the mean for all $|\tilde{z}|\gtrsim 1$. This indicates that while the potential fluctuations are extremely dynamically important near the midplane (where they can be key drivers of the evolution of vertical structure in the stellar disk), they may be negligible beyond one scale height away from it.
@@
Overplotted as black dashed curves in all panels are the predictions of the model, given by Section IV.1 for the mean and Equation 54 for the fluctuations. As expected given the excellent fit to the mean vertical profile shown in Figure 11, the model matches the measured mean vertical potential quite accurately. For each $\tilde{k}$ value of the fluctuations, the model curve typically falls within the $\pm 1\sigma$ range around the measured time-average curve, although for the smallest $\tilde{k}$ values, the power is somewhat overpredicted at $|\tilde{z}|\gtrsim 1$, and for the larger $\tilde{k}$ values, the power is somewhat underpredicted for $|\tilde{z}|\lesssim 1$. These discrepancies can be attributed to the fact that the underlying gas volume density fluctuations are actually more concentrated near the midplane than the mean vertical density profile. In the stellar-dynamical context, this difference may be consequential for vertical scattering processes experienced by stars in the thin disk, with small vertical excursions from the midplane compared to the gas scale height. Therefore, we describe an extension to our model allowing for the fluctuation vertical profile to differ from the mean in Appendix C. Nevertheless, we find that even our simple model with a single function $\zeta_{\mathrm{m}}$ for both the $k=0$ and $k>0$ components captures many of the essential features of the spectrum of potential fluctuations.
@@
Figure 13: The square root of the power spectrum of potential fluctuations as a function of $\tilde{k}$ at fixed $\tilde{z}=\{0,1,1.5\}$ (increasing from top to bottom in each panel) for the R8 (left, purple) and LGR4 (right, orange) simulations. We normalize by $\phi_{0}$; see Section IV.1. This is the same quantity shown in panels (c) and (d) of Figure 12, except now as a function of $\tilde{k}$ instead of $\tilde{z}$. For clarity, we only show the standard deviation (shaded region) around the time-averaged quantity (solid curve) for one value of $\tilde{k}$, but the fractional variations are comparable across all the curves. In each panel, the black dashed curve shows the result assuming a separable density profile with the fitted vertical profile shown in Figure 11, according to Equation 54. The dotted curves indicate power laws of slope indicated by the text in each panel for reference.
@@
Figure 13 shows the same root-mean-square potential fluctuation amplitudes as the bottom row of Figure 12, but this time as a function of $\tilde{k}$ for the midplane $\tilde{z}=0$ (top curve in each panel), as well as $\tilde{z}=1$ and $\tilde{z}=1.5$ (middle and bottom curves respectively). Once again, we show the $\pm 1\sigma$ range about the time-average quantity for just one curve for clarity, but the fractional variations about the mean are comparable for all $\tilde{z}$ values. We again overplot the predictions from the model given by Equation 54 as black dashed curves. We indicate the approximate power law scalings measured from the simulation at low $\tilde{k}$ on each panel and plot dotted lines of the corresponding slope for comparison.
[…]
@@
Thus far, we have been comparing our model to the simulation outputs by studying only the amplitudes of each Fourier component of the fluctuations. However, another consequence of Equation 49 is that the phase of $(\delta\phi_{\mathbf{k}})_{\mathrm{m}}$ is entirely set by the phase of $\delta_{\mathbf{k}}$, since all the other factors are real. One simple way to assess whether the separable model is adequate in this regard is to reconstruct a potential fluctuation field from the measured $\delta_{\mathbf{k}}$ at a given snapshot (by summing up the components given by Equation 49 using Section II) and comparing to the actual field at that time. If the phases of $\delta_{\mathbf{k}}$ and $\delta\phi_{\mathbf{k}}$ are well-correlated, then the reconstruction at each value of $z$ should have a similar morphology to the real field, but if the phases are not well-correlated, it may look entirely different.
@@
Ultimately, we have demonstrated that the simple separable model described above gives a good description of the potential fluctations as a function of both space and time. Together with the results regarding the linear surface density fluctuations presented in Section III, Equation 54 provides a direct input for stellar dynamical studies of e.g., diffusion processes driven by external perturbations (J. Binney & C. Lacey, 1988), which we will explore in a future work (Modak et al. in prep).
V Discussion
In this work, we have characterized the statistics of the (2D) surface density fluctuation fields of the ISM, and subsequently demonstrated that the structures of the (3D) volume density and gravitational potential fields are well-approximated by a separable model in which the 3D fluctuation fields essentially inherit the statistical structure of the 2D field (see Equation 38 and Equation 39), with a specified vertical profile (see Equation 44). Thus, our results can be summarized using our characterization of the surface density fluctuation spatio-temporal spectra in Section III.3. With a realization of the surface density fluctuation $\delta(\mathbf{x},t)\equiv\Sigma(\mathbf{x},t)/\overline{\Sigma}(t)-1$ in hand, realizations of the corresponding volume density, vertical gravitational potential, and gravitational potential fluctuations can then be produced using Equation 38, Section IV.1, and Equation 49 respectively, and center-of-mass motion (Equation 36) or thickness variation (Equation 37) can be specified as desired, calibrated to the values reported in Table 4.
Physically, the content of Section III.3 is that
@@ V Discussion
    fluctuation power is distributed across these modes as a steep power law with index $n_{\delta}\approx-2.2$ for linear surface density fluctuations $\delta$ (or $n_{s}\approx-2.8$ for logarithmic surface density fluctuations $s\equiv\ln(1+\delta)$), so that a majority of the power is concentrated at the largest spatial scales.
Below, we speculate on the implications and possible origins of the power-law spatial structure (Section V.1), discuss how our results may be compared with observations (Section V.2), and finally comment on caveats and possible generalizations of our characterization (Section V.3).
V.1 Implications and Origins of Power-Law Spectra
An immediate consequence of the power-law spatial spectra of the surface density field is that ISM substructures are not well-described as compact “clouds” with well-defined sizes and masses. In contrast to these classical models, the spectra we measure are relatively featureless, suggestive of a fractal, hierarchical structure. As an example of an application where this difference leads to qualitatively dissimilar results, we note that in the stellar-dynamical context, where it is natural to classify fluctuations by their scale (e.g., C. Hamilton et al. 2026), the large-scale ISM fluctuations that carry the most of the power drive scattering processes distinct from those of small-scale clouds. We will describe the implications of this fluctuation structure for stellar orbital heating and migration processes in Galactic disks in a future work (Modak et al. in prep).
@@ V.1 Implications and Origins of Power-Law Spectra
[…]
In the discussion above, we have explicitly written $P_{\mathrm{3D}}$ to denote power spectra measured in 3D (and averaged over angles for $k=(k_{x}^{2}+k_{y}^{2}+k_{z}^{2})^{1/2}$). For an isotropic spectrum, integrating vertically to obtain the 2D projection simply selects $k_{z}=0$, with $P(k_{x},k_{y})\propto P_{\mathrm{3D}}(k=(k_{x}^{2}+k_{y}^{2})^{1/2})$ up to a normalization factor. The spectra we find, $P_{\delta}\propto k^{-2.3}$ to $k^{-2.2}$, are considerably flatter than would be expected if we were simply projecting from the isotropic, single-phase models described above. The difference likely has more to do with the multiphase character of the gas than the fact that we are considering turbulence in a disk, since the ratio between the vertical dimension of the disk and the horizontal dimension in the simulation is modest, $2H/L\sim 0.6$. However, the cold gas is collected in small clouds with densities two orders of magnitude larger than those of the warm diffuse gas, while the density in superbubbles that fill a large fraction of the domain is two or more orders of magnitude below that of the warm gas.
V.2 Comparison with Observations